# No-go for the formation of heavy mass Primordial Black Holes in Single Field Inflation

### v5

@@ IX.1 Details of the one loop computation of two point function using in-in formalism in USR phase
It is important to note that, in the SR and USR regions, the second slow roll parameter behaves as a constant for which one can approximately consider . However, this approximation does not hold good at the conformal time scales at and . So instead of having the contribution from the full time scale the only significant contributions will appear at and . Consequently, the associated conformal time integral part as appearing in the above mentioned correlation functions can be evaluated by considering the following approximated expression:
@@
IX.2 Computation of cut-off regulated one-loop momentum integrals
@@
IX.2.1 One-loop momentum integrals in USR period
@@
where we define a new momentum and conformal time dependent function , which is defined by the following expression:
Here the Bogoliubov coefficients and in the USR region is given by:
After substituting the specific mathematical form of the momentum and conformal time dependent function in equation (110), we found:
where the four contributions , can be written as:
The outcomes of the aforementioned integrals are as follows after a few algebraic operations:
Let’s put the contributions from equations (119), (120),(121) and (122)) together now to better understand how the findings behave in the USR regime, which will result in the following results that have been simplified:
Then considering both the contribution from the short range as well as the long range quantum fluctuations we found the following simplified result:
To obtain this simplified result we have neglected all other suppressed contributions from the oscillatory terms as well as inverse power law terms of , as all of them will not be able to significant change the overall behaviour of the scalar power spectrum in USR period. The obtained results suggests that in the one-loop result IR sensitive logarithmic divergent and quadratic UV divergent contributions are dominating over all the other terms appearing in this computation. With the detailed analysis we have shown in the text portion of the paper that the quadratic UV divergence can be completely removed by implementing the well known adiabatic renormalization scheme. Additionally, by implementing the DRG resummation technique we have shown that the IR sensitivity can be further softened and shifted to the next order of perturbation theory. It’s also crucial to note that in order to extract the finite contributions from the one-loop momentum integrals in the USR period, we have constrained the momentum integration within a window, , by introducing two well known physical cut-offs: the IR cut-off and the UV cut-off which are the corresponding cut-off scales of UV and IR divergences inserted through the implementation of cut-of regularization before implementing the adiabatic renormalization and DRG resummation.
IX.2.2 One-loop momentum integral in SR period
@@
where we define a new momentum and conformal time dependent function in presence of Bunch Davies quantum initial condition, which is defined as:
After substituting the explicit mathematical form of the above function in the equation (127), we get the following simplified result:
Then considering both the contribution from the short range as well as the long range quantum fluctuations we found the following simplified result:
Here in SR phase represents the pivot scale which is expected to be in the present framework under consideration. In the SR region the final one-loop result will be controlled by the above mentioned IR logarithmic divergent contribution and a quadratic UV divergent contribution. We have shown in this paper by applying adiabatic renormalization scheme UV divergence can be completely by choosing the underlying renormalization scale properly. Also we have shown that IR divergence can be softened and shifted to next order in the perturbation theory.

### v6

@@ Appendix A Details of the one loop computation of two point function using in-in formalism in USR phase
It is important to note that, in the SR and USR regions, the second slow roll parameter behaves as a constant for which one can approximately consider . However, this approximation does not hold good at the conformal time scales at and . So instead of having the contribution from the full time scale the only significant contributions will appear at and . Consequently, the associated conformal time integral part as appearing in the above mentioned correlation functions can be evaluated by considering the following approximated expression [51]:
@@
Appendix B First step on renormalization
Translating the computation into a more comprehensible language will allow us to link it to the traditional renormalization techniques employed in quantum field theory with more ease. In this part, instead of utilizing several methods of this kind, we will just focus on the process of eliminating divergences at the level of the unrenormalized/bare action through the introduction of counter-terms. Ultimately, this will provide the renormalized form of the action, where, upon successful completion of the technique, any potentially harmful divergences—specifically, quadratic UV divergence—may be entirely eliminated and logarithmic IR divergences smoothed.
Let us start with the following bare action:
where the bare coupling parameters are defined as:
The renormalized form of the third order action for the comoving scalar curvature perturbation can be obtained by using the rescaling ansatz of the gauge invariant modes, which is very helpful in figuring out the relationship between the renormalized, unrenormalized/bare, and counter-term contributions. The renormalized version of the third-order action is given by the following expression:
Renormalized contributions are denoted by the subscripts R, B, and C, respectively. Firstly, it should be noted that the amount , or more accurately , is often called the counter-term. To estimate this amount directly, we must use the renormalization condition.
The expression for the third-order unrenormalized/bare action must then be translated into terms of the renormalized version using the previously mentioned rescaled renormalized form of the curvature perturbation. This may be simply accomplished by carrying out the following tasks:
The renormalized coupling parameters of the third-order perturbed action may be expressed in terms of the contributions of the bare and counter-terms as follows, using the previously given ansatz:
(113) -
In this case, the couplings and the independent operator contributions may be expressed as follows:
(114) In this instance, we have used the subsequent universal scaling relationship:
(115) In this case, the dotted contributions represent the higher-order terms in the associated power-series expansion. All higher-order minor effects have been disregarded, and we have limited our study to first-order variables. Consequently, this suggests that we have completed the last calculations needed to ascertain the exact contributions of the counter-terms within the linear domain of the parallel expansion.
Thus, for the gauge invariant comoving curvature perturbation, the third-ordered renormalized version of the perturbed actions may be obtained by completing all the previously specified steps:
where , the bare component of the third-order perturbed action, is specified directly in the equation (110). The third-order action may be further explained by the following phrase once the counter-terms have been introduced:
In the one-loop corrected primordial power spectrum, the counter-term contribution of the third-order action that is capable of eliminating the quadratic UV divergence contribution entirely is represented by the following expression:
Likewise, the following expression characterizes the counter-term contribution of the third-order action that might soften the behavior of the logarithmic IR divergence contribution in the one-loop adjusted primordial power spectrum:
Utilizing the above-mentioned facts, for scalar modes, the complete regularized and renormalized one-loop corrected power spectrum may thus be written as follows:
where the definitions of IR and UV counter-terms are:
Specifically, the UV counter can be further expressed in terms counter-terms of the coupling parameters as appearing in the action as stated in equation (118) for the SR and USR phases are given by the following expressions:
Next, we use the renormalization condition to get the expression for the counter-terms as they show up in the derived result previously described. The current problem may be understood clearly using Renormalization Group (RG) flow, which will also successfully correct the structure of any counter-terms arising in the calculation that are UV and IR sensitive. The renormalized spectrum is represented in Fourier space, and this is made feasible by the flow equation and matching beta functions provided for the appropriate 1PI one-loop corrected renormalized two-point amplitude. In this case, the Callan–Symanzik equation for this cosmological arrangement may be expressed as follows:
It is now crucial to keep in mind that, in this case, the appropriate total differential operator may be further minimized using the following formula:
where we define:
This further reduces to the following differential equation:
where the following describes the beta functions:
It is possible to compute the flow equations below by identifying the IR and UV counter-terms in the current environment at the renormalization scale:
The renormalized form of the spectral tilt for the scalar modes is described by the following first flow equation:
(133) -
The renormalized form of the spectral tilt running for the scalar modes is described by the following second flow equation:
(134) -
The renormalized form of the running of the spectral tilt running for the scalar modes is described by the third flow equation, which is as follows:
The flow equations that come before it demonstrate that the two-point amplitude of the power spectrum of primordial scalar modes exhibits a scale-dependent feature when the IR and UV counter-term effects are seen in each individual expression. When we apply the renormalization criterion, the structure of the counter-terms that are both UV and IR-sensitive will be fixed. We will now proceed with this utilizing the established facts at the CMB pivot scale , which prevents us from taking into account the following important constraints, which are explained in terms of renormalization conditions:
Condition I: The first renormalization requirement is that the tree-level contribution recorded during the first SR phase must precisely match the two-point amplitude of the scalar power spectrum following renormalization at the CMB pivot scale . Technically speaking, this claim may be expressed as follows:
(136) -
Condition II: The second renormalization criteria is that the logarithmic derivative of the amplitude of the scalar power spectrum with respect to the momentum scale following renormalization must exactly match the tree-level contribution estimated in the first SR phase at the CMB pivot scale . In a technical sense, this claim may be stated as follows:
(137) -
Condition III: The third renormalization requirement, which is the second logarithmic derivative of the two-point amplitude of the scalar power spectrum with respect to the momentum scale, must be fully satisfied by the tree-level contribution computed during the first slow-roll phase. Alternatively, this claim can be expressed technically as:
(138) -
Condition IV: According to the fourth renormalization criterion, the tree-level contribution computed in the first slow-roll phase at the CMB pivot scale must exactly equal the third logarithmic derivative of the two-point amplitude of the scalar power spectrum with respect to the momentum scale. Stated otherwise, the technical form of this claim is as follows:
Further limits on the parameters of the UV and IR-sensitive counter-terms are obtained as a direct consequence of the four renormalization requirements previously mentioned. A detailed summary of these limitations is provided below:
Consequence I: The following constraint condition expresses the immediate result of the first renormalization condition:
(140) -
Consequence II: The subsequent constraint condition expresses the immediate result of the second renormalization requirement:
(141) -
Consequence III: The third renormalization condition’s immediate result is conveyed by the following constraint condition:
(142) -
Consequence IV: The following constraint condition expresses the immediate result of the fourth renormalization condition:
Our thorough analysis of the issue revealed that, in order to meet the previously acquired sets of constraint criteria, we must very carefully define the UV counter-term . Only if we appropriately eliminate the quadratic divergence component can this be accomplished. Upon completion, the IR counter-term will have its form automatically fixed. Notwithstanding the aforementioned limits at the CMB pivot scale, determining the precise form of the UV counter-term is incredibly challenging at this level. The primary challenge at the technical level arises from the necessity for counter-terms in the SR and USR phases to independently eliminate the quadratic divergences’ contribution. As simple as it may appear, completing the computation we have done so far is rather challenging from a technological standpoint. The significance of the next three sections lies in the way the adiabatic renormalization approach allows us to fully eliminate the quadratic UV divergences’ contributions from each of the SR and USR phases separately. Following completion of this, we may use the condition to quickly ascertain the explicit form of the IR counter-term. Together with the quadratic divergence-free result for that comes from the adiabatic renormalization scheme, we have used this condition in the context of power spectrum renormalization.
We firmly believe that providing a clear explanation of the relationship between adding a counter-term at the level of action—a standard approach within the framework of Quantum Field Theory—and the renormalization schemes for adiabatic/wave function, and power spectrum will aid in understanding the applicability of the results derived in this paper. To prevent any further misunderstandings regarding the renormalization schemes employed or the interconnection among various tools and techniques used in this paper, let us discuss it in detail in the following subsections.
In the current computation, the counter-term contribution of the third order perturbed action—which we have indicated in our computation by in the equation (118)—is directly linked to the complete removal of the harmful quadratic UV divergence. It is feasible to clearly demonstrate as an immediate result of the computation with this particular section that the sum of the counter-terms for the five operators that were previously described may be represented in terms of a cumulative factor. We determine this factor to be the counter-term contribution that, when applied to the wave function/adiabatic renormalization scheme—a topic we will address in the following portion of this paper—would entirely eliminate the quadratic UV divergence.
However, in our computation, we have denoted the third-order perturbed action as in the equation (119), which is directly linked to the coarse-graining and smoothing of the logarithmic IR divergence’s behaviour. It is feasible to clearly demonstrate that the single counter term represents the instantaneous result of the calculation using this particular portion. This factor will smooth the logarithmic IR divergence’s behaviour by moving it to the higher order during the computation of the 1PI one-loop corrected two-point amplitude. This higher order corresponds to the higher even loop diagrams that arise in the perturbative expansion. Later in this work, we shall address this issue in depth in the context of the power spectrum renormalization technique.
Further, we justify more deeply the applicability of two renormalization schemes within one theoretical framework that we are studying in this paper. At the level of third-order perturbed action representing the counter-term contributions we found that the UV and IR sensitive terms are completely decoupled, as presented in equation (118) and equation (119). On the other hand, in the expression for the renormalized power spectrum the UV and IR counter terms appear as a product. Just by applying the previously mentioned renormalization conditions, one cannot fix the structure of both of these counter-terms simultaneously. For this reason, the adiabatic renormalization scheme, which responds to the quadratic divergences successfully, is used to determine the UV counter-term. Implementation of this scheme helps to remove quadratic divergences fully from this computation. This fixed structure of the UV counter-term is inserted in the expression for the power spectrum and further using the renormalization conditions one can further fix the structure of the IR counter-term, which helps to smoothen the logarithmic divergences. The latter we identify as the power spectrum renormalization scheme. It seems like we have used two completely separate schemes for the renormalization of a single theory, however, from the discussions presented in this section and its underlying connections with the renormalization in Quantum Field Theory of curved space-time it is now quite clear that these two schemes of renormalizations, i.e. adiabatic and power-spectrum are related to each other. Without using the adiabatic scheme, one cannot further use the power-spectrum scheme, as IR counter-terms cannot be fixed without fixing UV counter-terms. In the next two subsections where the details of the adiabatic and wave function renormalization schemes are presented, we further establish these claims in more detail.
Appendix C Computation of cut-off regulated one-loop momentum integrals
@@
C.1 One-loop momentum integrals in USR period
@@
(144) where we define a new momentum and conformal time dependent function , which is defined by the following expression:
(145) Here the Bogoliubov coefficients and in the USR region is given by:
(146) (147) After substituting the specific mathematical form of the momentum and conformal time dependent function in equation (144), we found:
(148) where the four contributions , can be written as:
(149) (150) (151) (152) The outcomes of the aforementioned integrals are as follows after a few algebraic operations:
(153) (154) (155) (156) Let’s put the contributions from equations (153), (154),(155) and (156)) together now to better understand how the findings behave in the USR regime, which will result in the following results that have been simplified:
(159) Here the symbol "Ei” signifies the exponential integral function, which is defined by the following expression:
(160) Then considering both the contribution from the short range as well as the long range quantum fluctuations we found the following simplified result:
(161) To obtain this simplified result we have neglected all other suppressed contributions from the oscillatory terms as well as inverse power law terms of , as all of them will not be able to significant change the overall behaviour of the scalar power spectrum in USR period. The obtained results suggests that in the one-loop result IR sensitive logarithmic divergent and quadratic UV divergent contributions are dominating over all the other terms appearing in this computation. With the detailed analysis we have shown in the text portion of the paper that the quadratic UV divergence can be completely removed by implementing the well known adiabatic renormalization scheme. Additionally, by implementing the DRG resummation technique we have shown that the IR sensitivity can be further softened and shifted to the next order of perturbation theory. It’s also crucial to note that in order to extract the finite contributions from the one-loop momentum integrals in the USR period, we have constrained the momentum integration within a window, , by introducing two well known physical cut-offs: the IR cut-off and the UV cut-off which are the corresponding cut-off scales of UV and IR divergences inserted through the implementation of cut-of regularization before implementing the adiabatic renormalization and DRG resummation.
C.2 One-loop momentum integral in SR period
@@
(162) where we define a new momentum and conformal time dependent function in presence of Bunch Davies quantum initial condition, which is defined as:
(163) After substituting the explicit mathematical form of the above function in the equation (162), we get the following simplified result:
(164) Then considering both the contribution from the short range as well as the long range quantum fluctuations we found the following simplified result:
(165) Here in SR phase represents the pivot scale which is expected to be in the present framework under consideration. In the SR region the final one-loop result will be controlled by the above mentioned IR logarithmic divergent contribution and a quadratic UV divergent contribution. We have shown in this paper by applying adiabatic renormalization scheme UV divergence can be completely by choosing the underlying renormalization scale properly. Also we have shown that IR divergence can be softened and shifted to next order in the perturbation theory.